# DiffArtist: Towards Structure and Appearance Controllable Image Stylization

### v3

@@ Abstract.
Artistic style includes both structural and appearance elements. Existing neural stylization techniques primarily focus on transferring appearance features such as color and texture, often neglecting the equally crucial aspect of structural stylization. In this paper, we present a comprehensive study on the simultaneous stylization of structure and appearance of 2D images. Specifically, we introduce DiffArtist, which, to the best of our knowledge, is the first stylization method to allow for dual controllability over structure and appearance. Our key insight is to represent structure and appearance as separate diffusion processes to achieve complete disentanglement without requiring any training, thereby endowing users with unprecedented controllability for both components. The evaluation of stylization of both appearance and structure, however, remains challenging as it necessitates semantic understanding. To this end, we further propose a Multimodal LLM-based style evaluator, which better aligns with human preferences than metrics lacking semantic understanding. With this powerful evaluator, we conduct extensive analysis, demonstrating that DiffArtist achieves superior style fidelity, editability, and structure-appearance disentanglement. These merits make DiffArtist a highly versatile solution for creative applications. Project homepage: https://github.com/songrise/Artist.
Generative art, Text-driven stylization, Structure and appearance, Stylization evaluation, Multimodal LLM applications,
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Arxiv; 2025; NYisbn: 978-1-4503-XXXX-X/2018/06submissionid: ccs: Computing methodologies Image processingccs: Applied computing Fine artsccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Figure 1. DiffArtist enables disentangled and fine-grained control of style strength from two orthogonal perspectives: structure and appearance. The prompt is “The Dream, by Picasso.”
@@
The essence of an artistic style lies not only in its visual appearance but also the structural elements (Gombrich and Gombrich, 1995; Jung et al., 2024). For example, the fragmentation of figures in Picasso’s Cubist works and the undulating sky in Van Gogh’s “Starry Night”, each contributing distinctly to their artistic expression. To automate the generation of artworks, existing neural stylization approaches (Huang and Belongie, 2017; Deng et al., 2022; Kwon and Ye, 2022; Wang et al., 2022; Huang et al., 2024; Cui et al., 2023) have predominantly focused on manipulating appearance-level attributes such as color and textures. The structural elements in the source image, however, are often modeled as the “content” and are explicitly preserved (Wang et al., 2024b; Wen et al., 2023; Yang et al., 2022). Consequently, these methods typically lack the capacity to curate meaningful structural-styles, substantially limiting their artistic potential and customizability.
Unlike appearance, structural stylization is inherently much more challenging as it demands artistically meaningful variations that (1) are aligned with the target style, (2) harmoniously integrated with the original image, and (3) without violating the semantic integrity of the source content. These objectives are at a higher, potentially semantics-level than the appearance-style, and, therefore, are less readily achieved. While descriptors like AdaIN (Huang and Belongie, 2017) and Gram loss (Gatys et al., 2015) may suffice for appearance-style modeling, the lack of adequate structure representation and a structure-style evaluator represent major obstacles in the development of structural stylization techniques. This challenge becomes even more pronounced in contemporary multimodal generation scenarios, where the style as textual prompt lacks a direct visual counterpart (Huang et al., 2024; Wang et al., 2023a; Cui et al., 2023; Hertz et al., 2024).
In search of the dual controllability of structure and appearance elements, the recent development of image generation models, in particular the diffusion models (DMs) has provided a promising solution. Compared with traditional stylization, the generative process allows for the sampling of stylized images of greater diversity in both structure and appearance. Building upon DMs, the stylization problem can re-framed as controlled image generation, where the reference image and style are additional conditions for sampling of stylized images. However, achieving this control in an artistically meaningful way is far from straightforward, as the diffusion process inherently entangles the generation of structure and appearance, as analyzed in this paper. This entanglement presents a major challenge for full controllability, resulting in a tradeoff between structural and appearance style strength (i.e., the S-A tradeoff). Consequently, existing diffusion-based approaches are either prone to undesired content modification (Chung et al., 2024; Lyu et al., 2023; Wang et al., 2024b) or suffer from constrained appearance-style strength (Wang et al., 2023b). Achieving dual controllability in the stylization thus remains an open question.
In this paper, we present the first comprehensive study for disentangled structure and appearance control for 2D stylization. We first present an styler, DiffArtist, which, to the best of our knowledge, is the first to enable the dual controllability for text-driven stylization. DiffArtist is based on explicit disentanglement of the structural and appearance generation in the diffusion process. It is plug-in with U-Net-based DMs, requiring neither optimization nor additional adapters like ControlNet (Zhang et al., 2023a) or IP-Adapter (Ye et al., 2023) (i.e., zero-shot), making it highly efficient for user personalization. This superior controllability overcomes the S-A tradeoff in DMs. Compared with conventional methods and other DM-based methods (e.g, (Wang et al., 2024a; Lin et al., 2024; Zhang et al., 2023a)), we demonstrate DiffArtist achieves significantly better dual controllability without deterioration in semantics, as exemplified in Fig. 1. Aside from structural stylization, it also allows appearance stylization with un-represented high fidelity, when configured with strict structure preservation, thereby unleashing full customization power.
In addition to the proposed styler, we further explore the development of art evaluators with human-aligned perceptions of structure and appearance style. The challenges for this exploration are threefold: (1) Structural style components are at a higher level than appearance and cannot be evaluated in isolation without considering their integration with content. (2) Evaluators must possess contextual awareness to detect structural artifacts that compromise semantic coherence. (3) Evaluators should also effectively associate textual style instructions with corresponding visual features for cross-modal evaluation. Multimodal LLMs (MLLMs), known for their robust zero-shot semantic-level understanding capabilities, emerge as an optimal solution to address these challenges. To this end, we propose (to the best of our knowledge) the first MLLM-based metrics for evaluating the structure and appearance in artistic stylization. We empirically demonstrate that this metric aligns significantly better with human artistic perception than existing metrics (Kirstain et al., 2023; Wang et al., 2004; Radford et al., 2021), paving the way for more human-aligned style evaluation. We summarize our contributions as follows:
@@
    We propose DiffArtist, the first stylization method that enables the dual controllability of structure and appearance-related style.
@@
    We present a novel MLLM-based evaluator for evaluation of structure and appearance in artistic stylization, which aligns better with human perception.
@@
    Extensive experiments demonstrate DiffArtist achieves superior stylization fidelity, with highly editable and disentangled control.
@@ 2. Related Works
The Stylization of Structure and/or Appearance. Structure and appearance collectively define the style of a visual representation. Existing neural stylization methods (Huang and Belongie, 2017; Deng et al., 2022; Kwon and Ye, 2022; Wang et al., 2022; Huang et al., 2024; Cui et al., 2023) has predominately focused on appearance stylization, generally achieved via encoder-decoder architecture. Only a few papers (Kim et al., 2020; Zhou et al., 2024) focus on transferring structural style components between 2D images. This can be achieved via image deformation. For instance, one may calculate a correspondence between the structure of content and style image, to perform a non-rigid deformation. However, these methods usually operate images in specific domains like portrait, requiring a stylized image as reference. They also do not allow for dual controllability. The dual controllability has recently been explored in image editing literature (Lin et al., 2024), yet they focus on targeted editing of realistic photos at single control level, instead of harmonious transfer between photo and artworks with high editablity. The structure stylization is more widely explored in 3D stylization, which has an explicit formulation of shapes (e.g, vertices (Michel et al., 2022), density (Jung et al., 2024)) and appearance (e.g., texture (Chen et al., 2023b) map, radiance (Wang et al., 2023a)), yet they cannot be directly applied to the 2D case. In this paper, we explore the first dual controllable 2D image stylization method.
@@
Text-Driven Image Stylization. Text-driven image stylization aims to stylize a source image according to style prompts. Early methods achieve it by optimizing certain image representation (Michel et al., 2022; Kwon and Ye, 2022; Wang et al., 2023a; Gal et al., 2022; Kim et al., 2022) with a multimodal similarity loss, usually implemented as the CLIP loss (Radford et al., 2021). Recently, it was discovered that text-to-image (T2I) DMs could also be adapted for similar optimization schemes (Poole et al., 2022; Hertz et al., 2023; Jiang et al., 2023; Kawar et al., 2023). These optimization-based methods are costly and slow, motivating recent methods explore the feed-forward paradigm. Instruct-Pix2Pix (Brooks et al., 2023) tune the diffusion model with a language model for generalized editing tasks. Diffstyler (Huang et al., 2024) learn a content and style-specific denoiser for disentanglement. FreeStyle (He et al., 2024) modulate U-Net feature for training-free stylization. Concurrent with this exploration is the stylized image generation. While related, they focus a different setting where the style is extracted from an image and the content is a prompt (Hertz et al., 2024; Wang et al., 2024a, b; Gao et al., 2024; Chen et al., 2023a). In this work, we focus on the structure and appearance control in text-driven stylization scenarios.
@@
Quantitative Evaluation of Style Transfer. Traditional metrics such as Gram Loss (Gatys et al., 2016), LPIPS (Zhang et al., 2018), and Fréchet Inception Distance (FID) (Heusel et al., 2017) evaluate quality of stylization from various perspectives. These metrics are at a relatively low-level, and are found to be misaligned with the holistic human artistic perception (Bo et al., 2018; So, 2023; Ioannou and Maddock, 2024; Chen et al., 2024; Yi et al., 2023; Wright and Ommer, 2022). To build art-specific evaluators, recent methods like ArtFID (Wright and Ommer, 2022) and ArtScore (Chen et al., 2024) attempt to quantify ”artness” of images, yet these methods cannot handle the open-vocabulary text-driven stylization, and they do not evaluate structure and appearance separately. In this paper, we propose a semantic-level MLLM-based evaluator to assess the structure and appearance fidelity, which aligns better with human perception.
@@ 3. Methodology
Figure 2. Overview of DiffArtist. We disentangle the structure and appearance of image as separate diffusion processes. All branches share the same U-Net parameters, and perform full denoising of $\tau$ steps. The feature pyramid from structure delegation is injected into the main branch for structure control, and the semantic-aware appearance is controlled via AdaIN.
3.1. Objective: Dual Controllability
Given a content image $I$, a style prompt $y$, and a U-Net-based DM $\mathcal{G}$, our objective is to control $\mathcal{G}$ to derive a distribution of stylized images $\hat{I}\sim\mathcal{G}(I,y)$ that harmoniously compose the provided content and style. Specifically, we aim to achieve disentangled controllability in both structure and appearance-related style. Our definition of structure and appearance in 2D imagery is mainly based on fine-art (Gombrich and Gombrich, 1995). A formal definition of them relates to visual semiotics (Van Leeuwen and Jewitt, 2000) and extend beyond the scope of this paper. Generally speaking, structure corresponds to the geometry, like contours and curvatures, while appearance corresponds to patterns like strokes and color pallette. We also aim to develop an evaluator $\mathcal{E}$, which can evaluate the fidelity of structure and appearance style in a way aligned with human perception.
The remaining parts of this section are organized as below. In Sec. 3.2, we review the basics of inversion-based image manipulation. In Sec. 3.33.4 we explain the motivation and design of control at a high-level, and details are described in Sec. 3.5. Sec. 3.6 outlines the proposed MLLM-based structure and appearance evaluators.
@@ 3.2. Preliminary: DDIM Inversion
To stylize a real image $\mathbf{x}_{0}:=I$ using DMs, inversion-based methods first approximate the noise latent $\mathbf{\tilde{x}}_{1:T}$ of it, typically using techniques such as DDIM inversion (Song et al., 2020). With the inverted latents, stylization can be achieved through generation with altered conditions (usually given as a style prompt $y$). Specifically, one may start with an intermediate step $\mathbf{x}_{\tau}$ (i.e., control point), where $\tau\in[1,T]$ for iterative DDIM sampling. Each denoising step is defined as:
@@
where $t$ is the timestep, and $\alpha_{1:T}$ is a specific noise schedule. The assumption in this paradigm is that with a proper $\tau$, the resulting stylized image $\hat{I}:=\hat{\mathbf{x}}_{0}$ harmoniously integrate the content of original image ${\mathbf{x}_{0}}$ with the style in prompt $y$.
@@ 3.3. Structure and Appearance in Noise Space
Existing neural stylization methods generally view an image as a composition of certain “content” and “style” components  (Li et al., 2017; Huang and Belongie, 2017; Wang et al., 2023b; Kotovenko et al., 2019; Sanakoyeu et al., 2018; Zhang et al., 2023b, 2020), where the style usually refers to the low-level visual features such as appearance. As this paper motivates to stylize both structure and appearance, we adopt a different modeling to decompose image as appearance and structure components: $\mathbf{x}_{0}=\mathcal{G}_{0}(\mathbf{z}^{a}_{0},\mathbf{z}^{s}_{0})$, where $\mathcal{G}_{0}=(\cdot,\cdot)$ is a composition function, $\mathbf{z}^{a}_{0}$ and $\mathbf{z}^{s}_{0}$ are the latent appearance and structure factorization, respectively. This is a “static” perspective. In the diffusion process, the distribution of image $x_{0}$ is tied with the intermediate distributions in $\mathbf{x}_{1:T}$, where a denoiser $\epsilon_{\theta}$ learns the transition $q_{\theta}(x_{t-1}|x_{t},y)$ via $\epsilon$-prediction. We therefore posit similar factorization of noise-residuals $\epsilon_{\theta}$:
@@
(2) $\epsilon_{\theta}(\mathbf{x}_{t},t;y)=\mathcal{G}_{t}\left(\kappa_{t},\psi_{t}\right),$
@@
where $\kappa_{t}$ and $\psi_{t}$ denote the structure and appearance representation at diffusion time-step $t$ (detailed later), and $\mathcal{G}_{t}(\cdot,\cdot)$ is the noise-space composition function at time $t$.
3.4. Structure and Appearance as Delegate Process
Controlling the structure and appearance-level stylization in DMs presents unique challenges due to inherent entanglement. We provide both theoretical and empirical evidence for this phenomenon in Appendix. B, revealing the technical challenge of S-A tradeoff in DM. The key insight from this analysis is that solidary control point $\mathbf{x}_{\tau}$ used in existing DMs-based stylization does not suffice for S-A disentanglement. Consequently, we argue that the structure and appearance shall be represented and controlled over complete and separate diffusion trajectories.

### v4

@@ Abstract.
Artistic styles are defined by both their structural and appearance elements. Existing neural stylization techniques primarily focus on transferring appearance-level features such as color and texture, often neglecting the equally crucial aspect of structural stylization. To address this gap, we introduce DiffArtist, the first 2D stylization method to offer fine-grained, simultaneous control over both structure and appearance style strength. This dual controllability is achieved by representing structure and appearance generation as separate diffusion processes, necessitating no further tuning or additional adapters. To properly evaluate this new capability of dual stylization, we further propose a Multimodal LLM-based stylization evaluator that aligns significantly better with human preferences than existing metrics. Extensive analysis shows that DiffArtist achieves superior style fidelity and dual-controllability compared to state-of-the-art methods. Its text-driven, training-free design and unprecedented dual controllability make it a powerful and interactive tool for various creative applications. Project homepage: https://diffusionartist.github.io.
Generative art; Text-driven stylization; Structure and appearance; Stylization evaluation; Multimodal LLM applications
copyright: acmlicensedjournalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755010isbn: 979-8-4007-2035-2/2025/10submissionid: $c$s: Computing methodologies Appearance and texture representationsccs: Computing methodologies Image manipulationccs: Applied computing Fine arts
@@ 1. Introduction
Figure 1. DiffArtist enables disentangled and fine-grained control of style strength from two orthogonal perspectives: structure and appearance. The style prompt is “The Dream, by Picasso.”
@@
The essence of an artistic style lies not only in its appearance—color and texture—but also its structure—geometry and composition (Gombrich and Gombrich, 1995; Jung et al., 2024; Goodman, 1976). For example, the fragmentation of figures in Picasso’s Cubist works and the undulating sky in Van Gogh’s “Starry Night”, each contributing distinctly to their artistic expression. Existing neural stylization approaches (Huang and Belongie, 2017; Deng et al., 2022; Kwon and Ye, 2022; Wang et al., 2022; Huang et al., 2024; Cui et al., 2023) predominantly focus on manipulating appearance-level attributes. The structural elements in the source image, however, are often viewed as part of “content” and are explicitly preserved (Wang et al., 2024b; Wen et al., 2023; Yang et al., 2022, 2023). This fundamental limitation prevents them from capturing the true essence of an art style, severely restricting their expressive potential and customizability.
The root of this limitation lies in the inherent complexity of structural stylization. Unlike appearance-style transfer, structural stylization requires a delicate balance between three competing objectives: (1) aligning with the target style, (2) harmonizing with the source image’s composition, and (3) preserving the core semantic integrity of the content. These objectives operate at a high-level semantic plane, exposing a critical gap in current methods. While descriptors like AdaIN (Huang and Belongie, 2017) and Gram loss (Gatys et al., 2015) may suffice for appearance-style modeling, the lack of adequate structure representation and structure-style evaluators presents significant obstacles in the development of structural stylization techniques. This challenge is amplified in recent multimodal generation scenarios, where a style prompt offers no explicit visual template (Huang et al., 2024; Wang et al., 2023a; Cui et al., 2023; Hertz et al., 2024).
The advent of Diffusion Models (DMs) offers a powerful new paradigm for achieving this dual controllability, as their generative sampling process enables far greater structural and appearance diversity than prior methods. This reframes stylization as a conditional generation task, guided by a source image and a style prompt (image or text). However, this generative power comes with a critical, unaddressed challenge: the diffusion process inherently entangles the generation of structure and appearance. We identify this as a fundamental Structure-Appearance (S-A) Tradeoff: intensifying structural changes inadvertently corrupts appearance style, while strengthening appearance washes out structural transformations. This tradeoff directly explains the core failures of existing diffusion-based methods, which are either prone to severe content degradation (Chung et al., 2024; Lyu et al., 2023; Wang et al., 2024b) or suffer from weak, constrained stylization (Wang et al., 2023b; Yang et al., 2023). Achieving dual controllability in the stylization thus remains an open question.
To solve this, we introduce DiffArtist, the first framework to our knowledge that offers explicit, disentangled control over both structure and appearance in 2D stylization. At its core, DiffArtist explicitly disentangles the structural and appearance generation as separated diffusion processes, with shared semantic information. This design directly overcomes the fundamental S-A tradeoff and functions as a zero-shot, plug-and-play module for any pretrained U-Net-based DM, requiring no costly fine-tuning or external adapters (Zhang et al., 2023a; Ye et al., 2023). As evidenced in this paper, this design provides true disentanglement—a key advantage over ControlNet-based methods (Wang et al., 2024a; Lin et al., 2024) where adjusting one style factor adversely impacts the other. As demonstrated in Fig. 1, this unprecedented level of control allows DiffArtist to achieve strong, semantically coherent stylization, unleashing the full creative potential of dual-style customization.
Evaluating this novel capability of dual control requires understanding on image semantics, where existing evaluation metrics obsolete. This exposes a critical need for a new evaluation paradigm. To address this, we introduce our second major contribution: a Multimodal LLM (MLLM)-based evaluator designed for dual stylization. We argue that any such evaluator must satisfy three key criteria: (1) operate at a high-level semantic plane to assess structure, (2) possess contextual awareness to maintain semantic integrity, and (3) perform robust cross-modal association between text prompts and visual forms. By leveraging the zero-shot reasoning of MLLMs, our proposed metric meets these criteria. We empirically show that it aligns significantly better with human artistic judgment than existing stylization metrics (Kirstain et al., 2023; Wang et al., 2004; Radford et al., 2021), establishing a more reliable and human-centric standard for future stylization research.
We summarize our contributions as follows:
@@
    We identify the S-A tradeoff in diffusion models as the key challenge for disentangled dual controllability.
@@
    We propose DiffArtist, the first 2D stylization method that enables the dual controllability of structure and appearance.
@@
    We present a novel MLLM-based evaluator for evaluating structure and appearance in artistic stylization, which aligns better with human perception.
  4. (4)
    Extensive experiments demonstrate that DiffArtist achieves superior stylization fidelity, control editability, and disentanglement than existing approaches.
@@ 2. Related Works
Figure 2. Overview of DiffArtist.Our method disentangles stylization by processing structure and appearance through two independent diffusion trajectories (delegations). At each denoising step, the main stylization branch is conditioned on semantic-level features from the structure and appearance delegations. All three branches share the same pretrained U-Net parameters, and perform full denoising of $\tau$ steps. The entire framework operates without requiring any fine-tuning or adapters.
@@
The Stylization of Structure and/or Appearance. Structure and appearance collectively define the style of a visual representation. Existing neural stylization methods (Huang and Belongie, 2017; Deng et al., 2022; Kwon and Ye, 2022; Wang et al., 2022; Huang et al., 2024; Cui et al., 2023) have predominantly focused on appearance stylization, generally achieved via an encoder-decoder architecture. Only a few papers (Kim et al., 2020; Zhou et al., 2024) focus on transferring structural style components between 2D images. This is usually accomplished by calculating a correspondence between content and style image and performing a non-rigid deformation. However, these methods typically operate on images in specific domains like portraits, requiring an in-domain stylized reference. Moreover, they do not enable dual controllability. Very recently, dual controllability has been explored in Ctrl-X (Lin et al., 2024), but its framework is designed for localized, image edits rather than the global, harmonious style transformations. Meanwhile, the concept of disentangled control is explored in 3D synthesis, where explicit representations of shape (e.g., meshes (Michel et al., 2022) and radiance fields (Wang et al., 2023a)) and appearance (e.g., texture mapping (Chen et al., 2023b)) make separation natural. The success in 3D, however, is not directly transferable to the 2D domain due to the lack of explicit geometric priors in single images. In this paper, we explore the first dual controllable 2D image stylization method.
@@
Text-Driven Image Stylization. Text-driven image stylization aims to stylize a source image according to style prompts. Early methods achieve it by optimizing certain image representation (Michel et al., 2022; Kwon and Ye, 2022; Wang et al., 2023a; Gal et al., 2022; Kim et al., 2022) with a multimodal alignment objective, typically implemented as the CLIP loss (Radford et al., 2021). Recently, it was discovered that text-to-image (T2I) DMs could also be adapted for similar optimization schemes (Poole et al., 2022; Hertz et al., 2023; Jiang et al., 2023; Kawar et al., 2023). These optimization-based methods are costly and slow, motivating recent exploration of the feed-forward paradigm. Instruct-Pix2Pix (Brooks et al., 2023) tunes the diffusion model along with a language model for generalized editing tasks. Diffstyler (Huang et al., 2024) learn a content and style-specific denoiser for disentanglement. FreeStyle (He et al., 2024) modulates the U-Net feature for training-free stylization. Concurrent with this exploration is the stylized image generation (Hertz et al., 2024; Wang et al., 2024a, b; Gao et al., 2024; Chen et al., 2023a). While related, they focus on a different setting where the style is extracted from an image and the content is a prompt. In this work, we focus on the structure and appearance control in text-driven stylization scenarios.
@@
Quantitative Evaluation of Style Transfer. Quantitatively evaluating style transfer is a long-standing problem. Initial approaches repurposed low-level metrics, including Gram Loss (Gatys et al., 2016), LPIPS (Zhang et al., 2018), and FID (Heusel et al., 2017). However, recent literature found that these metrics are fundamentally incapable of capturing the holistic and semantic qualities of human artistic perception (Bo et al., 2018; So, 2023; Ioannou and Maddock, 2024; Chen et al., 2024; Yi et al., 2023; Wright and Ommer, 2022). In response to these shortcomings, art-specific evaluators like ArtFID (Wright and Ommer, 2022) and ArtScore (Chen et al., 2024) were developed to better quantify the abstract concept of ”artness”. Nevertheless, they cannot handle open-vocabulary text-driven stylization and lack the mechanism to evaluate structure and appearance separately. This paper propose a semantic-level MLLM-based evaluator to assess the structure and appearance fidelity, which aligns better with human perception.
@@ 3. Methodology
3.1. Objective: Disentangled Dual Controllability
Given a source image $I$, a text-based style prompt $y$, and a pretrained diffusion model $\mathcal{G}(\cdot)$, our primary objective is to generate a stylized image $\hat{I}$ that preserves the semantic content of $I$ while harmoniously embodying the style described by $y$. The core innovation we pursue is disentangled dual control, meaning that decomposing the style prompt $y$ into two orthogonal components—structure and appearance—and controlling their strength independently. Our definition of structure and appearance in a 2D image is mainly based on fine art (Gombrich and Gombrich, 1995). A formal definition of them is challenging as it relates to visual semiotics (Van Leeuwen and Jewitt, 2000; Goodman, 1976), extending beyond the scope of this paper. Generally speaking, structure corresponds to the shapes, like contours and curvatures, while appearance corresponds to local patterns, like strokes and color palettes. We also aim to develop an evaluator $\mathcal{E}$, which can evaluate the fidelity of structure and appearance style in a way aligned with human perception.
The remaining parts of this section are organized as follows. In Sec. 3.2, we review the basics of inversion-based image manipulation. In Sec. 3.33.4, we explain the motivation and design of control at a high level, and details are described in Sec. 3.5. Sec. 3.6 outlines the proposed MLLM-based structure and appearance evaluators.
@@ 3.2. Preliminary: DDIM Inversion
To stylize a source image $\mathbf{x}_{0}:=I$ using DMs, inversion-based methods first approximate the noise latents $\mathbf{x}_{1:T}$ of $I$, achieved via techniques such as DDIM inversion (Song et al., 2020). Stylization is then performed through re-generation with altered conditions (usually specified as a style prompt $y$). Specifically, one may start with an intermediate step $\mathbf{x}_{\tau}$ (i.e., control point), where $\tau\in[1,T]$ for iterative DDIM sampling. Each denoising step is formulated as follows:
@@
where $\epsilon_{\theta}$ is the denoiser, $t$ is the timestep, and $\alpha_{1:T}$ is a predefined noise schedule. The assumption of this paradigm is that with a proper $\tau$, the resulting stylized image $\hat{I}:=\hat{\mathbf{x}}_{0}$ harmoniously integrates the structure and appearance of style in prompt $y$ with the source image $I$.
@@ 3.3. Structure and Appearance in Noise Space
Prevailing neural stylization methods are built on a paradigm that separates an image into “content” and “style” (Li et al., 2017; Huang and Belongie, 2017; Wang et al., 2023b; Kotovenko et al., 2019; Sanakoyeu et al., 2018; Zhang et al., 2023b, 2020; Ding et al., 2024). In this view, the style usually refers to the feature maps extracted from certain layers of a neural network. To advance stylization towards both structure and appearance controllability, we adopt different modeling that decomposes an image as its structure and appearance components: $\mathbf{x}_{0}=\mathcal{G}_{0}(\mathbf{z}^{s}_{0},\mathbf{z}^{a}_{0})$, where $\mathcal{G}_{0}=(\cdot,\cdot)$ is a composition function, $\mathbf{z}^{s}_{0}$ and $\mathbf{z}^{a}_{0}$ are the latent structure and appearance factorization, respectively. This is a “static”, image-level perspective. In the diffusion process, the distribution of image $x_{0}$ is tied with the intermediate distributions in $\mathbf{x}_{1:T}$, where the denoiser $\epsilon_{\theta}$ learns the transition $q_{\theta}(x_{t-1}|x_{t},y)$ via $\epsilon$-prediction. Therefore, we posit similar factorization of predicted noise residual, which is a “dynamic” decomposition across the full frequency bands:
@@
(2) $\epsilon_{\theta}(\mathbf{x}_{t},t;y)=\mathcal{G}_{t}\left(\kappa_{t},\psi_{t}\right),\quad t\in[0,T]$
@@
where $\kappa_{t}$ and $\psi_{t}$ denote the structure and appearance representation at diffusion time-step $t$ (detailed later), respectively. $\mathcal{G}_{t}(\cdot,\cdot)$ is a conceptual noise-space composition function at time $t$.
3.4. Structure and Appearance as Delegate Diffusion Process
We argue that the fundamental obstacle to achieving dual control in diffusion-based stylization is the inherent entanglement of structure and appearance. Our analysis, detailed in Appendix D, pinpoints the source of this problem: the reliance on a single latent trajectory $\mathbf{x}_{\tau}\xrightarrow{}\mathbf{x}_{0}$. This monolithic generation process forces structural and appearance attributes to compete for influence at every denoising step, creating the S-A tradeoff that fundamentally limits controllability.